OpenStudy (michaelbp):

John, Johan, and Jonathan collect postage stamps. Johan and Jonathan have the same number of stamps. Two times Johan’s collection is eleven less than five times John’s collection, and three times Jonathan’s collection is three less than seven times John’s collection. The number of postage stamps in John’s collection is

OpenStudy (chrishogan):

Hello, I can help you. :)

OpenStudy (michaelbp):

ok

OpenStudy (chrishogan):

Let d = number of stamps from John (4 letters, d is 4th letter, etc) Let e = number of stamps from Johan Let h = number of stamps from Jonathan Johan and Jonathan have the same number of stamps, so: e = h 2 times Johan's collection is 11 less than 5 times John's collection, so: 2e = 5d - 11 3 times Jonathan's collection is 3 less than 7 times John's collection,so: 3h = 7d - 3 Now we have three equations and three unknowns. You're asked for the number of stamps in John's collection, so you want the value to variable "d". Since e = h, we can substitute "h" for "e" in the second equation. 2h = 5d - 11 3h = 7d - 3 now we have two equations and two unknowns. I'll use elimination, starting with multiplying the first equation by -3 and the second one by 2. -6h = -15d + 33 6h = 14d - 6 now add both together and you have an equation with one unknown remaining: 0 = -d + 27 d = 27 So your answer is 27 stamps.

OpenStudy (chrishogan):

Fan me so I can help you with anymore questions you might have in the future.

OpenStudy (chrishogan):

Also make sure to close your question after it's been answered. :)